Prove the Postulate of Pasch
1
Expert's answer
2020-10-12T15:50:23-0400

Pasch's Postulate states that if a line intersects a triangle not at a vertex, then the line intersects two sides of the triangle.

i.e given a "\begin {aligned}
\triangle
\end {aligned}" ABC and a line segment "\begin {aligned}
l
\end {aligned}", if "\begin {aligned}
l
\end {aligned}" intersects the segment AB, it also intersects the segment AC or BC.



"\underline\bold{Proof}"


Note:

  1. "l" divides plane "ABC" into two distinct planes
  2. Call the two planes and "H"1 and "H"2
  3. Point A and Point B are in different half planes since "l" divides line AB
  4. Let A "\varepsilon" plane "H"1 and B "\varepsilon" plane "H"2
  5. C "\varepsilon" "H"1 or C "\varepsilon" "H"2

Case 1

Assuming Point C "\varepsilon" "H"1

Since Points A and C "\varepsilon" "H"1, "\therefore" A and C are on the same side of the plane

This also means that "l" does not intercept AC

Since C "\varepsilon" "H"1 and B "\varepsilon" "H"2, "\therefore" B and C are on opposite sides of "l"

"\therefore" "l" intercepts BC


Case 2

Assuming Point C "\varepsilon" "H"2

Since Points B and C "\varepsilon" "H"2, "\therefore" B and C are on the same side of the plane

This also means that "l" does not intercept BC

Since C "\varepsilon" "H"2 and A "\varepsilon" "H"1, "\therefore" A and C are on opposite sides of "l"

"\therefore" "l" intercepts AC



Therefore "l" intercepts either AC or BC